OpenStudy (anonymous):

Angle C is a right angle in triangle ABC. The measure of angle B is 54 degrees. The length of AB is 15 units. Find the measure of angles A and the length of sides AC and BC. Round your answer to two decimal places.

OpenStudy (anonymous):

This is what I got: 180 - 54 - 90 = 36degrees tan(36) = x/15 x = 10.89 (15^2) + (10.89^2) = 18.53 AC = 18.53 BC = 10.89

OpenStudy (anonymous):

u know the answer?

OpenStudy (anonymous):

looks right to me :)

OpenStudy (anonymous):

im not sure whether it is right.

OpenStudy (anonymous):

hmmm strange. i only received partial credit. anything that could be wrong?

OpenStudy (anonymous):

i think it is right

OpenStudy (anonymous):

oh i see what went wrong. AB is the hypotenuse. its the longest side. Theres no way you can have a side longer than that.

OpenStudy (anonymous):

i'll draw a picture and post it. When you did the question did you draw a picture of it or just do it in your head?

OpenStudy (anonymous):

i tried to do it in my head. to no avail. haha. a picture would be great.

OpenStudy (anonymous):

with these type of problems, you definitely want to draw a quick picture, just so you make sure everything is straight. If i had a nickel for everytime i got a problem wrong doing it in my head, i would have $175.45 >.>

OpenStudy (anonymous):

hahahhaha thanks a lot. much appreciated